OpenStudy (anonymous):

How to sovle this for "x": xt^2=x-2 Please explain it step by step thanks ...

OpenStudy (anonymous):

is this \[xt^2=x-2\]?

OpenStudy (mr.math):

\[xt^2-x=-2 \implies x(t^2-1)=-2 \implies \dots\]

OpenStudy (anonymous):

Thanks man :)

OpenStudy (mr.math):

You're welcome! :)